OpenStudy (faman39):

I have a question, how do you reduce large fraction fast? If you don't remember more multiplication.

OpenStudy (welshfella):

find their prime factors

OpenStudy (faman39):

for example 456/678

OpenStudy (welshfella):

456 = 2 * 2 * 2* 3 * 19 678 = 2 * 3 * 113 cancel out the duplicates in this case they are 2 and 3 so we have 2* 2* 19 / 113 = 76/113

OpenStudy (welshfella):

that took me a while - then I'm an old guy! lol

OpenStudy (faman39):

Do we usually need to reduce all large fraction with 2 or 3 ? yeah, I was thinking that how we can reduce large fraction like that faster if we don't remember longest or bigger multiplications

OpenStudy (welshfella):

I don't think there's a faster way of doing this for large numbers All the division you do are pretty simple. You can use a calculator of course.

OpenStudy (mathmate):

@Faman39 It depends on whether you can divide mentally by one-digit numbers.

OpenStudy (faman39):

Actually, I need to take tabe test within few month. They dont allow us to use calculator and they give really short time to solve those type of problem.

OpenStudy (welshfella):